solve a system step-by-step of linear equations using Gaussian elimination.

2x + 7y + z = 1

x + 3y - z = 2

x + 7y + 12z = 45



1
Expert's answer
2021-03-09T03:55:07-0500
"\begin{cases}
 2x + 7y + z = 1 \\
 x + 3y - z = 2 \\
 x + 7y + 12z = 45 \\
\end{cases}"

"\begin{pmatrix}
 2 & 7 & 1 &| 1 \\
 1 & 3 &-1 & |2 \\
 1 & 7 & 12 &|45 \\
\end{pmatrix} \underrightarrow{R_3 + (-R_2)}
\begin{pmatrix}
 2 & 7 & 1 &| 1 \\
 1 & 3 &-1 & |2 \\
 0 & 4 & 13 &|43 \\
\end{pmatrix} \underrightarrow{R_2 + (-\frac{1}{2}R_1)}
\begin{pmatrix}
 2 & 7 & 1 &| 1 \\
 0 & -\frac{1}{2} &-\frac{3}{2} & |\frac{3}{2} \\
 0 & 4 & 13 &|43 \\
\end{pmatrix} \\"


"\\
\\
\underrightarrow{R_3 + 8R_2}
\begin{pmatrix}
 2 & 7 & 1 &| 1 \\
 0 & -\frac{1}{2} &-\frac{3}{2} & |\frac{3}{2} \\
 0 & 0 & 1 &|55 \\
\end{pmatrix} \underrightarrow{\frac{1}{2}R_1 ; -2R_2;}
\begin{pmatrix}
 1 & \frac{7}{2} & \frac{1}{2} &| \frac{1}{2} \\
 0 & 1 & 3 & | -3 \\
 0 & 0 & 1 &|55 \\
\end{pmatrix} \Longrightarrow"


"\Longrightarrow 
\begin{cases}
 x + \frac{7}{2}y + \frac{1}{2}z = \frac{1}{2} \\
 y + 3z = -3 \\
 z = 55 \\
\end{cases}""z = 55;\\ y + 3 \cdot 55 = -3; \\ y = -168; \\
x + \frac {7}{2} \cdot (-168) + \frac{1}{2} \cdot 55 = \frac{1}{2}; \\ x -588 + 27,5 = 0,5; \\
x= 561;"

Answer: "x = 561; y = -168; z = 55."